Circular No. 6039

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

AG DRACONIS R. Viotti, Istituto Astrofisica Spaziale, Consiglio Nazionale delle Ri-

cerche, Frascati; R. González-Riestra, International Ultraviolet Explorer (IUE) Observatory, European Space Agency, Madrid; and C. Rossi, Istituto Astronomico, La Sapienza University, Rome, write: “We observed the symbiotic star AG Dra with the IUE during its present outburst phase (IAUC 6009) on June 29 and July 4, 5, and 12. The ultraviolet continuum fluxes at 134.0 and 286.0 nm are 4.1 and 2.9 × 10^{−12} erg cm^{−2} s^{−1} A^{−1} on July 12, respectively, whereas one week earlier they were lower by factors 1.25 and 1.10. The strongest emission line is He II at 164 nm, with an observed intensity of 1.7 × 10^{−10} erg cm^{−2} s^{−1} on July 5 (25 percent stronger than during the 1981 outburst). At low resolution, there are also intense emissions of N V, C IV, the O III Bowen lines, the He II Paschen lines, and the O IV] + Si IV 140-nm blend. At high resolution, the He II 164-nm line shows a narrow emission core and extended wings, with a FWHM of 0.4 nm. The ultraviolet spectrum is qualitatively similar to that of the 1981 outburst, but line and continuum fluxes appear more intense.”

LQ SAGITTARII U. Munari and T. Zwitter, Asiago Astrophysical Observatory; and

A. Bragaglia, Astronomical Observatory, Bologna, communicate: “A lowresolution, optical/near-infrared spectrogram of LQ Sgr (= N Sgr 1897), recently announced as undergoing a symbiotic-nova outburst (cf. IAUC 6018), was secured on July 14.1 UT with the Boller & Chivens spectrograph of the 1.5-m European Southern Observatory telescope. The spectrum is typical of a Mira variable close to maximum brightness. It does not resemble that of any known symbiotic or classical nova at maximum or declining phases. We therefore suggest that LQ Sgr is a normal Mira variable.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) H. C. Bhatt, Indian Institute of Astrophysics, Bangalore, reports that

he detected an intense flash in the infrared H (1.65-μm) band corresponding to the impact of the fragment S = 5 of P/Shoemaker-Levy 9 on Jupiter, with an InSb infrared photometer on the 0.75-m telescope at the Vainu Bappu Observatory, Kavalur. The brightness began to increase sharply, with a rise time of ∼ 2 s at July 21.63843 UT. The flash lasted 32 s with a peak flash brightness of ∼ −2.7 mag in the H band.

1994 July 27 Daniel W. E. Green